A solution was prepared by adding 12.00 ml 3.00 mL
Answer #1

1. From the given data

a. concentration [I-] = 0.13 m x 12 ml/36 ml = 0.043 M

b. concentration [S2O8^2-] = 0.13 m x 15 ml/36 ml = 0.054 M

c. concentration [S2O3^2-] = 0.13 M x 5 ml/36 ml = 0.018 M

d. from the above calculations,

ln[I-] = -3.15

e. from the above calculations,

ln[S2O8^2-] = -2.92

f. rate = -(-2.92/109) = 0.027 M/s

2. with rate = k[A]^2[B]

with given concentration of [A] and [B] and initial rate value we get,

rate constant k,

k = rate/[A]^2[B]

   = 5.08 x 10^-7/(8.72 x 10^-3)^2.(4.93 x 10^-2)

   = 0.1355 M-2.s-1